1. Show (graphically) how MC pricing can result in losses for the monopoly.
Answer #1

If price is set such that P=MC,then

TR=P*Q=15*8=120

TC=AC*Q=19*8=152

Since TC>TR,the monopoly is incurring losses from this approach.